1.] the equation of a locus is x2+y2-4x+6y-3=0.Show that the equation becomes x2+y2=16 if the origin is shifted.

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Ashwathy

x2+y2-4x+6y-3=0

(x-2)2 + (y+3)2 =16


this is a equation of circle having center ( 2,-3)

so now shift the origin at (2,-3)

x' =x+2

y'=y-3

so equation become x2+y2 =16

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin